# Hierarchical Dataset Selection for High-Quality Data Sharing

### v1

@@ Abstract
https://plan-lab.github.io/projects/dash
@@ 1 Introduction
To solve this new task, we propose Dataset Selection via Hierarchies (DaSH), a hierarchical Bayesian method that models dataset utility at both the group and dataset levels. Given a large pool of candidate datasets, grouped based on dataset origin (e.g., institution or collection) DaSH performs structured exploration to infer both group-level relevance and individual dataset utility via posterior inference over observed model performance. This hierarchical modeling allows DaSH to prioritize informative groups and avoid wasted evaluation on unrelated or harmful sources. Experiments on two benchmarks demonstrate DaSH significantly outperforms state-of-the-art baselines by up to 26.2% in accuracy under low-resource settings. The contributions of this work are:
@@
    We benchmark DaSH against four state-of-the-art data selection methods across two public datasets, demonstrating consistent performance improvements, improves accuracy by up to 26.2% Digit-Five and 10.8% on DomainNet. Ablation studies show DaSH remains robust to grouping noise and scales effectively to large dataset pools, whereas existing methods frequently select irrelevant or low-utility data samples.
@@ 2 Related Work
Hierarchical Bandits. Hierarchical bandit algorithms address decision-making problems where actions are structured in a hierarchy, enabling efficient exploration and exploitation across multiple levels (hong2022hierarchical; munos2014bandits). In recommendation systems, hierarchical bandits have been employed to model user preferences (yue2012hierarchical) and item categories (wang2018online; zuo2022hierarchical), enabling personalized content delivery under resource constraints through adaptive frameworks (yang2020hierarchical; santana2020contextual). Beyond recommendation, hierarchical bandits have been applied to intelligent tutoring, decentralized reinforcement learning, and multi-task off-policy learning (castleman2024hierarchical; hong2023multi; kao2022decentralized). These applications highlight the flexibility of hierarchical formulations in structuring complex decision processes across domains. Concurrently, theoretical advancements have focused on regret minimization and generalization across tasks using hierarchical Bayesian models (kveton2021meta; hong2022hierarchical; guan2024improved), offering principled frameworks for exploration under structured priors. Inspired by works in this space, our method tackles the unique setting of dataset selection by introducing a hierarchical Bayesian formulation that propagates dataset utility estimates across groups, enabling efficient amortization of training feedback via structured priors, and improving robustness to irrelevant or redundant datasets. To our knowledge, this is the first work to employ hierarchical bandits to dataset selection, with empirical evidence showing large gains in both accuracy and efficiency over non-hierarchical alternatives.
@@ 3 Method
where $\mu_{i}$ is the mean of the prior distribution for data group $g_{i}$, $\sigma_{i}^{2}$ is the variance of the group prior, $\hat{\sigma}_{i}^{2}$ is the variance of the dataset prior $\theta_{i,j}$, and $\sigma_{r}^{2}$ is the variance of the reward observation model. The goal is to iteratively update the posterior distribution of $\theta_{i}$ and $\theta_{i,j}$ by incorporating all observed reward values accumulated up to the current time step $t$. Through this continual update process, DaSH converges towards accurate estimations of the true distributions for both $\theta_{i}$ and $\theta_{i,j}$ after a number of iterations, as described Algorithm 1 in Appendix. Initialization begins with all dataset groups $\mathbf{g}$ sharing a common prior $\mathcal{N}(\mu_{0},\sigma_{0}^{2})$ and $\mathcal{N}(\theta_{0},\hat{\sigma}_{0}^{2})$. At each time step $t$, $\hat{\theta}_{i}$ is drawn from the normal distributions associated with each dataset group $\hat{\theta}_{i}\sim P(\theta_{i}|r_{i})$ and the dataset group $g_{i}$ with the largest value is chosen. Given dataset group selection $g_{i}$, DaSH then draws $\hat{\theta}_{i,j}$ from the distributions associated with the datasets within the chosen dataset group, i.e., $\hat{\theta}_{i,j}\sim P(\theta_{i,j}|r_{i,j})$, and selects the dataset with the largest values, denoted as $D(t)\!=\!d_{i,j}$.
@@ 4 Experiments
Figure 3 summarizes the empirical obtained by training local models on different external datasets. These ground-truth results serve as a reference for evaluating the potential benefit of dataset selection. In Digit-Five, models trained on external datasets consistently underperform compared to their local baselines, indicating strong domain-specific bias. In contrast, DomainNet exhibits more favorable cross-domain transfer; for example, training the REAL classifier on subsets from CLIPART yields noticeable performance gains. This distinction underscores the practical relevance of dataset selection in heterogeneous sharing scenarios.
Baselines.
Core-sets (sener2018active), which selects representative samples via geometric coverage, such that model learned only on the selected subset are as competitive.
@@ 5 Ablation Studies
We evaluate the ability of each method to identify useful datasets under stringent exploration budgets. Specifically, each method explores each dataset only once, totaling 15 steps across the 15 datasets in Digit-Five. Figure 5 reports the resulting accuracy for each domain. Under this extreme budget constraint, both DaSH and DaSH (mixed) outperform the non-hierarchical DaS (flat) in 4 out of 5 domains. The gains over DaS (flat) are substantial: +5.2% on MNIST, +6.0% on SVHN, +7.4% on USPS, and +9.0% on SYN. Even with imperfect grouping, DaSH (mixed) closely tracks the performance of perfect grouping, with accuracy differences within 1–2% in most domains. The Local and Global baselines show that hierarchical variants close more than half the gap to the global optimum despite operating under a 15-step budget. These results confirm that hierarchical grouping enables efficient, high-quality dataset selection even under severe exploration limits.
@@ 7 Conclusion
This work addresses a key bottleneck in machine learning: selecting training datasets from diverse sources such as institutions, repositories, or collections. We introduce DaSH, a dataset selection framework that models the hierarchical relationship among datasets and data sources to improve selection efficiency and downstream performance. Experimental results demonstrate that DaSH consistently outperforms non-hierarchical and existing instance-level data selection baselines, and remains robust under realistic constraints such as imperfect grouping and limited exploration budgets. These findings underscore the importance of effectively automating practical data curation as machine learning models increasingly depend on large-scale heterogeneous data sources from various online repositories. Future directions include incorporating multi-objective selection criteria such as utility, fairness, and domain coverage and applying DaSH to large-scale, multi-institutional data sharing platforms, where group membership and dataset availability evolve over time.
8 Acknowledgments

### v2

@@ Abstract
Project Pagehttps://plan-lab.github.io/projects/dash
@@ 1 Introduction
To solve this new task, we propose Dataset Selection via Hierarchies (DaSH), a hierarchical Bayesian method that models dataset utility at both the group and dataset levels. Given a large pool of candidate datasets, grouped based on dataset origin (e.g., institution or collection), DaSH performs structured exploration to infer both group-level relevance and individual dataset utility via posterior inference over observed model performance. This hierarchical modeling allows DaSH to prioritize informative groups and avoid wasted evaluation on unrelated or harmful sources. Experiments on two benchmarks demonstrate DaSH significantly outperforms state-of-the-art baselines by up to 26.2% in accuracy under low-resource settings. The contributions of this work are:
@@
    We benchmark DaSH against four state-of-the-art data selection methods across two public datasets, demonstrating consistent performance gains, improving accuracy by up to 26.2% Digit-Five and 10.8% on DomainNet. Ablation studies show DaSH remains robust to grouping noise and scales effectively to large dataset pools, whereas existing methods frequently select irrelevant or low-utility data samples.
@@ 2 Related Work
Hierarchical Bandits. Hierarchical bandit algorithms address decision-making problems where actions are structured in a hierarchy, enabling efficient exploration and exploitation across multiple levels (hong2022hierarchical; munos2014bandits). In recommendation systems, hierarchical bandits have been employed to model user preferences (yue2012hierarchical) and item categories (wang2018online; zuo2022hierarchical), enabling personalized content delivery under resource constraints through adaptive frameworks (yang2020hierarchical; santana2020contextual). Beyond recommendation, hierarchical bandits have been applied to intelligent tutoring, decentralized reinforcement learning, and multi-task off-policy learning (castleman2024hierarchical; hong2023multi; kao2022decentralized). These applications highlight the flexibility of hierarchical formulations in structuring complex decision processes across domains. Concurrently, theoretical advancements have focused on regret minimization and generalization across tasks using hierarchical Bayesian models (kveton2021meta; hong2022hierarchical; guan2024improved), offering principled frameworks for exploration under structured priors. Inspired by works in this space, our method tackles the unique setting of dataset selection by introducing a hierarchical Bayesian formulation that propagates dataset utility estimates across groups, enabling efficient amortization of training feedback via structured priors, and improving robustness to irrelevant or redundant datasets. To our knowledge, this is the first work to employ hierarchical bandits for dataset selection, with empirical evidence showing large gains in both accuracy and efficiency over non-hierarchical alternatives.
@@ 3 Method
where $\mu_{i}$ is the mean of the prior distribution for data group $g_{i}$, $\sigma_{i}^{2}$ is the variance of the group prior, $\hat{\sigma}_{i}^{2}$ is the variance of the dataset prior $\theta_{i,j}$, and $\sigma_{r}^{2}$ is the variance of the reward observation model. The goal is to iteratively update the posterior distribution of $\theta_{i}$ and $\theta_{i,j}$ by incorporating all observed reward values accumulated up to the current time step $t$. Through this continual update process, DaSH converges towards accurate estimations of the true distributions for both $\theta_{i}$ and $\theta_{i,j}$ after a number of iterations, as described in Algorithm 1 in the Appendix. Initialization begins with all dataset groups $\mathbf{g}$ sharing a common prior $\mathcal{N}(\mu_{0},\sigma_{0}^{2})$ and $\mathcal{N}(\theta_{0},\hat{\sigma}_{0}^{2})$. At each time step $t$, $\hat{\theta}_{i}$ is drawn from the normal distributions associated with each dataset group $\hat{\theta}_{i}\sim P(\theta_{i}|r_{i})$ and the dataset group $g_{i}$ with the largest value is chosen. Given dataset group selection $g_{i}$, DaSH then draws $\hat{\theta}_{i,j}$ from the distributions associated with the datasets within the chosen dataset group, i.e., $\hat{\theta}_{i,j}\sim P(\theta_{i,j}|r_{i,j})$, and selects the dataset with the largest values, denoted as $D(t)\!=\!d_{i,j}$.
@@ 4 Experiments
Figure 3 summarizes the empirical results obtained by training local models on different external datasets. These ground-truth results serve as a reference for evaluating the potential benefit of dataset selection. In Digit-Five, models trained on external datasets consistently underperform compared to their local baselines, indicating strong domain-specific bias. In contrast, DomainNet exhibits more favorable cross-domain transfer; for example, training the REAL classifier on subsets from CLIPART yields noticeable performance gains. This distinction underscores the practical relevance of dataset selection in heterogeneous sharing scenarios.
@@
Baselines.
Core-sets (sener2018active), which selects representative samples via geometric coverage, such that models learned only on the selected subset are as competitive.
@@ 5 Ablation Studies
We evaluate the ability of each method to identify useful datasets under stringent exploration budgets. Specifically, each method explores each dataset only once, totaling 15 steps across the 15 datasets in Digit-Five. Figure 5 reports the resulting accuracy for each domain. Under this extreme budget constraint, both DaSH and DaSH (mixed) outperform the non-hierarchical DaS (flat) in 4 out of 5 domains. The gains over DaS (flat) are substantial: +8.8% on MNIST, +1.8% on USPS, +9.8% on MNIST-M, and +4.5% on SYN. Even with imperfect grouping, DaSH (mixed) closely tracks the performance of perfect grouping, with accuracy differences within 1–2% in most domains. The Local and Global baselines show that hierarchical variants close more than half the gap to the global optimum despite operating under a 15-step budget. These results confirm that hierarchical grouping enables efficient, high-quality dataset selection even under severe exploration limits.
@@ 7 Conclusion
This work addresses a key bottleneck in machine learning: selecting training datasets from diverse sources such as institutions, repositories, or collections. We introduce DaSH, a dataset selection framework that models the hierarchical relationship among datasets and data sources to improve selection efficiency and downstream performance. Experimental results demonstrate that DaSH consistently outperforms non-hierarchical and existing instance-level data selection baselines, and remains robust under realistic constraints such as imperfect grouping and limited exploration budgets. These findings underscore the importance of effectively automating practical data curation as machine learning models increasingly depend on large-scale heterogeneous data sources from various online repositories. Future directions include incorporating multi-objective selection criteria such as utility, fairness, and domain coverage, and applying DaSH to large-scale, multi-institutional data sharing platforms, where group membership and dataset availability evolve over time.
Acknowledgments